You need to borrow by taking a 12-year loan. Your budget allows us to pay $980 at the end of each year for the length of the loan’s term as the loan payment. Annual rate is 5%. How much can you borrow today?
Answer #1

Ans $ 8685.99

P = Regular Payments
PV = Loan Amount
r = rate of interest
n = no of periods
P = r (PV)
1 - (1 + r )^-n
980 = 5% *PV
1 - (1 / (1 + 5%)^12))
980 = 0.05 * PV
0.443162581822441
PV = 980 / 0.05 * 0.443162581822441
PV = 8685.99